# revtex4-2Repair the float

### v1

@@ Abstract
Quantum convolutional neural networks (QCNNs) are quantum circuits for recognizing quantum phases of matter at low sampling cost and have been designed for condensed matter systems in one dimension. Here we construct a QCNN that can perform phase recognition in two dimensions and correctly identify the phase transition from a Toric Code phase with $\mathbb{Z}_{2}$-topological order to the paramagnetic phase. The network also exhibits a noise threshold up to which the topological order is recognized. Our work generalizes phase recognition with QCNNs to higher spatial dimensions and intrinsic topological order, where exploration and characterization via classical numerics become challenging.
@@
The remarkable progress in quantum computing technologies has led to the development of noisy intermediate-scale quantum (NISQ) computers [1] as well as to proof-of-principle demonstrations of quantum error correction [2, 1], fault-tolerant gate sets [3] and algorithms [4]. As a consequence, the output quantum states of quantum computers and simulations are becoming too complex to characterize them with classical computers. While many local properties of quantum states can be efficiently characterized using shadow tomography [5], the characterization of global properties remains an open challenge.
The direct processing of quantum data on quantum computers offers a potential solution to this problem, where recent developments include quantum autoencoders [6, 7], quantum principle component analysis [8], certification of Hamiltonian dynamics [9, 10], and quantum reservoir processing [11]. Particularly promising are quantum neural networks that use parameterized quantum circuits, measurement and feed-forward to characterize large amounts of quantum data [12, 13, 14, 15, 16].
A central field for applications of quantum computing is condensed matter physics, where the characterization of quantum states is of central importance. In particular, the classification of phases of matter [17, 18] is a key requirement for understanding strongly correlated materials [19]. While phases of matter in one-dimensional systems are well-understood thanks to accurate classical numerical simulations [20], two- and higher-dimensional systems are notoriously hard to simulate on classical computers due to the exponential growth of Hilbert space with the number of particles [21, 22].
@@
Topological many-body phases of matter have been prepared on quantum computers using unitary quantum circuits [23] as well as measurement and feed-forward [24, 25], thus opening a new way for their investigation. The characterization of these topological phases however remains challenging due to the absence of local order parameters and because other characteristic quantities, such as fidelity susceptibility and entanglement, are hard to measure. Previous approaches to address this challenge include classical machine learning techniques applied to data from classical numerical simulations [26, 27, 28] and classical shadows [29]. Topological phases have also been detected from data measured in an experiment by combining techniques from quantum error correction and renormalization-group flow [30], which is a well-established method for the classification of phases of matter [19]. Nonetheless, these approaches are limited by the large amounts of data required for describing quantum states on classical computers.
@@
Quantum convolutional neural networks (QCNNs) have been proposed to substantially reduce the sample complexity of recognizing symmetry-protected topological phases [13, 31] in one dimension and were shown to uncover characteristic properties of such phases from training data via supervised learning [13, 32, 33]. There are explicitly constructed QCNNs that are robust against errors in quantum data [34] and therefore implementable on a superconducting quantum processor under NISQ conditions [35]. However, these QCNNs have only been applied to one-dimensional systems, which can be accurately simulated on classical numerical computers. The extension of QCNNs to two- and higher-dimensional systems, where quantum computers can potentially provide a computational advantage, has remained unexplored.
@@
Here, we address this open challenge by generalizing QCNNs to detect $\mathbb{Z}_{2}$-topological order in two-dimensional spin arrays. We analytically construct a QCNN inspired by quantum error correction and Multiscale Entanglement Renormalization Ansatz [36]. Using matrix-product-state simulations, we show that the QCNN recognizes the topological phase of the Toric Code Hamiltonian in a magnetic field from a topologically trivial magnetic phase. In contrast to methods based on the processing of measurement data on a classical computer [30], the QCNN precisely identifies the critical value of the magnetic field where the Toric Code undergoes the phase transition. Moreover, the QCNN output is robust against incoherent errors below a threshold error probability allowing for the application of the QCNN under NISQ conditions on current quantum computers.
QCNNs are phase recognition circuits, which are an analog of classical Convolutional Neural Networks (CNNs). A QCNN can be written as a unitary of the form [13, 35, 31, 34],
$U_{\mathrm{QCNN}}=U_{\mathrm{FC}}\prod_{j}^{d}U_{\mathrm{CP}}^{(j)}\,,$ (1)
where the unitary $U_{\mathrm{CP}}^{(j)}$ represents a combination of a convolutional and a pooling layer, $d$ is the number of subsequent layers and the unitary $U_{\mathrm{FC}}$ is a fully connected layer that prepares the output. Subsequently, the output qubit(s) are measured. For the $j^{\mathrm{th}}$ layer in the QCNN structure, we specifically define the unitary $U_{\mathrm{CP}}^{(j)}$ as $U_{\mathrm{CP}}^{(j)}=U_{\mathrm{P}}^{(j)}U_{\mathrm{C}}^{(j)}(U_{\mathrm{C}}^{(j-1)})^{\dagger}$, where $U_{\mathrm{P}}^{(j)}$ is the unitary implemented by the pooling procedure, $U_{\mathrm{C}}^{(j)}$ the convolution of the current layer and $(U_{\mathrm{C}}^{(j-1)})^{\dagger}$ the adjoint of the convolution of the preceding layer $j-1$.
If we construct the convolutional layers from Clifford gates and the pooling procedure via single qubit Pauli-operations that are controlled by the results of measurements of other qubits, we can shorten the QCNN circuit dramatically by pushing $U_{\mathrm{C}}^{(j)}$ through the pooling circuit $U_{\mathrm{P}}^{(j)}$ [35, 37]. We then obtain a modified pooling layer $\tilde{U}_{\mathrm{P}}^{(j)}$ and subsequently cancel all convolutions except the first one, such that we can write the QCNN unitary as
$U_{\mathrm{QCNN}}=U_{\mathrm{FC}}\left(\prod_{j}^{d}\tilde{U}_{\mathrm{P}}^{(j)}\right)U_{\mathrm{C}}^{(0)}\,.$ (2)
Our algorithm is based on the fact that topological order is robust to local perturbations [17]. Hence, one can define characteristic states |$\ket{\psi_{i}}$ for each phase of the Hilbert space and treat other states of the same phase as characteristic states masked by unitary perturbations |$\ket{\phi_{j}}=U_{j,\mathrm{E}}\ket{\psi_{j}}$ that we identify as errors. If one can recover the characteristic state by removing these errors, one has successfully identified the corresponding phase. In Figure 1 we sketch such a phase space and the recovery of the reference state by using error correction. The explicit structure of our QCNN can be found in the supplementary material.
@@
[width=1]standalone/RGflow1dim
Figure 1: Schematic of phase recognition via a QCNN for two phases with characteristic states |$\ket{\psi_{1}}$ and |$\ket{\psi_{2}}$. Other states |$\ket{\phi_{j}}$ in the respective phases can be generated by local unitary perturbations |$\ket{\phi_{j}}=U_{j,\mathrm{E}}\ket{\psi_{j}}$ and the phase can for each input state be identified by undoing the respective perturbation $U_{j,\mathrm{E}}$ via sufficient error correction.
At each qubit layer $l$, we can generate an output by ‘measuring’ the qubits in the computational basis,
$M^{l}_{\mathrm{QCNN}}=\dfrac{2}{N_{l}}\left(\sum_{j=1}^{N_{l}}m_{j}\right)-1\,$ (3)
where $m_{j}\in\{-1,1\}$ is the measurement outcome of qubit $j$ and $N_{l}$ the number of qubits in layer $l$. Specifically, we implement the controlled Pauli gates of the pooling circuit classically on the bit string of computational basis measurements for all qubits and update the bit string accordingly. As these operations are performed in classical post-processing, the ’measurements’ correspond to simply recording the classical data. For our QCNN, $M^{l}_{\mathrm{QCNN}}$ as defined in Equation 3, will have the following averages over a large number of samples
$\braket{\mathrm{QCNN}}\coloneqq\braket{M^{l}_{\mathrm{QCNN}}}=\begin{cases}1&\text{if topological phase}\\ 0&\text{if non-topological phase}\,.\end{cases}$ (4)
We consider the two-dimensional Toric Code [38] with periodic boundary conditions as a natural candidate for exploring QCNNs for two-dimensional systems as it shows intrinsic $\mathbb{Z}_{2}$-topological order [38]. As previous approaches to QCNNs have been restricted to one-dimensional settings [13, 35, 33, 37], the Toric Code lends itself ideally to expand the field of QCNNs to higher dimensional systems. The Toric Code is a two-dimensional square lattice with the qubits located on its edges. For exploring its condensed matter physics, it can be described by the Hamiltonian
$H_{\mathrm{TC}}=-\sum_{i}A_{i}-\sum_{i}B_{i}$ (5)
with the plaquette operators $A_{i}=\prod_{i=1}^{4}X_{i}$ and the star operators $B_{i}=\prod_{i=1}^{4}Z_{i}$, where $X_{i}$ ($Z_{i}$) is the Pauli-X(Z) operator on qubit $i$, see Figure 2.
@@
[width=]standalone/QCNN_diagram2
@@
Figure 2: QCNN structure. The first lattice on the input side shows the Toric Code lattice with dimension $d=4\times 4$, which is periodic in vertical and horizontal direction. The qubits on the respective sublattices are represented by gray (horizontal edges) and green circles (vertical edges). Examples of plaquette ($A_{i}$) and vertex ($B_{i}$) operators are highlighted in blue and orange. The circuit of the convolution maps the stabilizers of the input lattice to measurements on individual qubits of the respective colors, i.e. blue qubits carry the measurement of the plaquette stabilizer that is anchored on the qubits to their left and orange qubits carry the information of stabilizers to their right. This convolution evolves error patterns on the input state and brings them into a form that can be corrected by the local operations in the pooling if the input was in the topological phase. Iterated pooling procedures reduce the lattice dimension (green squares) with every step until only two output qubits are left (green dots) and read out.
@@
If the Toric Code is exposed to external magnetic fields of sufficient strength, it will undergo a second-order phase transition [39, 40] between the topologically ordered and a paramagnetic phase. To explore these regimes, we modify Equation 5 to include applied magnetic fields in X- and Z-direction with strengths $h_{X}$ and $h_{Z}$,
$H_{\mathrm{mag}}=H_{\mathrm{TC}}-h_{Z}\sum_{i}^{N}Z_{i}-h_{X}\sum_{i}^{N}X_{i}\,.$ (6)
@@
The concept of the QCNN that we apply to (noisy) ground states of this model is sketched in Figure 2 and its detailed description can be found in the supplementary material. The QCNN we here propose is capable of detecting the topologically ordered phase with high accuracy, that is, the QCNN’s output converges to 1 if the input state is in the topological phase and to 0 otherwise.
@@
To evaluate the performance of our QCNN we carried out several numerical simulations. For investigating the scaling of its performance with increasing depth $d$, we need to consider large system sizes, for which we use Matrix Product State (MPS) simulations [41] to approximate ground states of the Hamiltonian Equation 6 and classical simulations of Clifford gates on Pauli strings for exploring the effects of incoherent Pauli noise.
@@
We first test the QCNN on ground states of $H_{\textrm{mag}}$ with a magnetic field in Z-direction, $h_{X}=0$ in Equation 6. We show the details of our MPS implementation in the supplementary material. With our implementation, we can generate the phase diagram as shown in Figure 3, where we consider field strengths $0\leq h_{Z}\leq 1$. From iDMRG simulations for our cylindrical lattice, we determine the phase boundary at $h_{Z}^{*}=0.34$ as a peak in the second derivative $d^{2}E/dh_{Z}^{2}$ of the ground state energy with respect to $h_{Z}$, which we mark in the phase diagram. The plot shows that the QCNN output $\braket{M^{l}_{\mathrm{QCNN}}}$, c.f. Equation 3, shows a transition at this expected point up to the sample resolution. Additionally, the sharpness of the transition increases with the layer number in the sense that the absolute value of the gradient of $\braket{M^{l}_{\mathrm{QCNN}}}$ increases, which makes the transition in the output more step-like with an increasing number of pooling layers. This shows the successful recognition of the topological phase by the QCNN.
@@
Figure 3: QCNN output for a variation of the magnetic field strength $h_{Z}$, where $h_{X}=0$, and for different depths $d$ of the QCNN, corresponding to the number of pooling layers $U_{P,j}$. Each data point represents 2000 samples from the MPS representation of the corresponding ground state. The shaded region corresponds to the topologically ordered phase and the vertical line to critical magnetic field strength $h^{*}_{Z}=0.34$. The insert shows the output for different samples with higher resolution around the phase transition, which are calculated with a bond dimension of $\chi=1500$, whereas for the main plot, the bond dimension is $\chi=1250$. For all samples, we observe increased steepness at the phase transition in the QCNN output with increasing QCNN depth corresponding to greater precision in the phase recognition.
Next, we test the QCNN on ground states of $H_{\textrm{mag}}$ with magnetic field in both X- and Z-direction. First, we consider constant $h_{X}=0.1$ leading to a non-vanishing probability of error syndromes $B_{i}=-1$ on vertices and thus $M_{\mathrm{QCNN}}^{0}<1$, see Figure 4. Second, we consider the transition into the trivial phase driven by perturbations in both X- and Z-directions for $h_{X}=h_{Z}=h$ [42]. The QCNN output rapidly decreases with increasing depth $d$ above the multicritical point $h=0.35$, see Figure 4b. This provides clear evidence that the QCNN recognizes the phase transition also for these cases.
Figure 4: a) Output of the QCNN for the MPS samples under a variation of $h_{Z}$ in the presence of a weak magnetic field in X-direction with $h_{X}=0.1$ with a bond dimension of $\chi=1250$. The output is displayed for simultaneous pooling in the X(Z)-basis for the corresponding QCNN depth $d_{X}$($d_{Z}$). The phase transition for $h_{Z}$ is correctly recognized in the presence of a field in X-direction. Furthermore, the output of the pooling layers $d_{Z}$ quickly converges to unity due to the weak $h_{X}=0.1$. b) Detection of the multicritical point [42] in the presence of X- and Z-field with $h=h_{X}=h_{Z}$ for a bond dimension of $\chi=1500$. We take the product of the pooling in the X- and Z-basis for each QCNN layer to show the combined phase recognition along the multicritical line $h_{X}=h_{Z}$. A split plot can be found in the supplementary material in Figure 10.
Figure 5: QCNN output for incoherent Pauli noise perturbing the Toric Code ground state as in Equation 7. Via tracking the evolution of Pauli strings on the lattice under the convolution and the pooling layers we can classically simulate the QCNN output for large system sizes. For this plot, we calculate the results for about 9.6 Million qubits, which allows us to employ more pooling layers compared to the MPS samples. As for the case of the magnetic field, we find a transition in the QCNN output that becomes more pronounced with increasing depth of the QCNN. The red vertical line marks the error threshold of $p_{\mathrm{th}}=2.28\%$.
@@
We also test whether our QCNN is able to recognize the topological phase in the presence of incoherent noise. We do this by considering the Toric Code ground state perturbed by Pauli noise. Specifically, we apply random Pauli noise to each individual qubit in the form of the error channel
$\rho=\mathcal{E}(\ket{\psi}\bra{\psi})=\sum_{l=0}^{m}K_{l}\ket{\psi}\bra{\psi}K_{l}^{\dagger}\,,$ (7)
with $K_{l}\in\{\sqrt{p_{\mathds{1}}}\mathds{1},\,\sqrt{p_{X}}X,\,\sqrt{p_{Z}}Z\}^{\otimes N}$ to the input qubits. $p_{X}$($p_{Z}$) is the probability for a Pauli-X(Z) error on any single qubit and $p_{\mathds{1}}=1-p_{X}-p_{Z}$. A Pauli-Y error occurs as the product of a Pauli-X and a Pauli-Z error on the same qubit. As the QCNN only consists of Clifford and controlled Pauli gates, we can then subsequently track the evolution of the Pauli error strings on the lattice under the operations of the QCNN. It is sufficient to employ the following gate identities $X_{j}=H_{j}Z_{j}H_{j}$, $\textrm{CNOT}_{i,j}X_{i}=X_{i}X_{j}\textrm{CNOT}_{i,j}$ and $\textrm{CNOT}_{i,j}Z_{j}=Z_{i}Z_{j}\textrm{CNOT}_{i,j}$, where $H_{j}$ is the Hadamard gate on qubit $j$ and $\textrm{CNOT}_{i,j}$ applies a Pauli $X$ on qubit $j$ conditioned on qubit $i$.
We present results for the QCNN output, Equation 3, for cases where the input state is the reference state of the topological phase, i.e. a ground state of $H_{\mathrm{TC}}$ as in Equation 5 perturbed by Pauli-Z noise in Figure 5, which shows a transition in the QCNN output over the Pauli-Z error probability $p_{Z}$. The absolute value of the gradient at the transition point increases with the number of QCNN layers. This corresponds to the desired behavior of the QCNN as the confidence of the QCNN output increases with additional layers. In the supplementary material, we also show results for simultaneous Pauli-Z and Pauli-X noise.

### v2

# Quantum Convolutional Neural Network for Phase Recognition in Two Dimensions

@@ Abstract
Quantum convolutional neural networks (QCNNs) are quantum circuits for characterizing complex quantum states. They have been proposed for recognizing quantum phases of matter at low sampling cost and have been designed for condensed matter systems in one dimension. Here we construct a QCNN that can perform phase recognition in two dimensions and correctly identify the phase transition from a Toric Code phase with $\mathbb{Z}_{2}$-topological order to the paramagnetic phase. The network also exhibits a noise threshold up to which the topological order is recognized. Furthermore, it captures correlations between all stabilizer elements of the Toric Code, which cannot be accessed by direct measurements. This increases the threshold for errors leading to such correlations and allows for correctly identifying the topological phase in the presence of strong correlated errors. Our work generalizes phase recognition with QCNNs to higher spatial dimensions and intrinsic topological order, where exploration and characterization via classical numerics become challenging.
@@
With the remarkable progress in quantum computing technologies Arute et al. (2019); Krinner et al. (2022); Postler et al. (2022); Bluvstein et al. (2023), the output states of quantum computers and simulations are becoming too complex to characterize them with direct measurements and classical computing. Although many local properties of quantum states can be efficiently characterized using shadow tomography Huang et al. (2020), the characterization of global properties remains an open challenge. Direct processing of quantum data on quantum computers offers a potential solution to this problem, where recent developments include quantum autoencoders Romero et al. (2017); Bondarenko and Feldmann (2020), quantum principle component analysis Lloyd et al. (2014), certification of Hamiltonian dynamics Wiebe et al. (2014); Gentile et al. (2021), and quantum reservoir processing Ghosh et al. (2019). Particularly promising are quantum neural networks that use parameterized quantum circuits, measurement and feed-forward  [^1] Feed-forward refers to a quantum operation conditioned on the outcome of a projective mid-circuit measurement, typically used in quantum error correction. to characterize large amounts of quantum data Farhi and Neven (2018); Cong et al. (2019); Beer et al. (2020); Kottmann et al. (2021); Gong et al. (2023).
@@
A central field for applications of quantum computing is condensed matter physics, where the characterization of quantum states is of central importance. In particular, the classification of phases of matter Kitaev (2003); Savary and Balents (2016) is a key requirement for understanding strongly correlated materials Sachdev (1999). While phases of matter in one-dimensional systems are well understood thanks to accurate classical numerical simulations McCulloch (2008), two- and higher-dimensional systems are notoriously hard to simulate on classical computers Zheng et al. (2017); Hangleiter et al. (2020). Of particular interest in this context are topological many-body phases of matter, which have already been prepared on quantum computers Satzinger et al. (2021); Semeghini et al. (2021); Bluvstein et al. (2022); Iqbal et al. (2023), thus opening new avenues for their exploration. However, the characterization of these topological phases remains challenging because of the absence of local order parameters. Other characteristic quantities, such as fidelity susceptibility and entanglement, are hard to measure. Previous approaches addressing this challenge include applying classical machine learning techniques to data from classical numerical simulations van Nieuwenburg et al. (2017); Carrasquilla and Melko (2017); Greplova et al. (2020) and classical shadows Huang et al. (2021). These approaches are however limited by the large amounts of data required for describing quantum states on classical computers.
@@
Quantum convolutional neural networks (QCNNs) have been proposed to substantially reduce the sample complexity of recognizing symmetry-protected topological phases in one dimension Cong et al. (2019); Lake et al. (2022) and were shown to uncover characteristic properties of such phases from training data via supervised learning Cong et al. (2019); Caro et al. (2022); Liu et al. (2022). Moreover, these QCNNs can be explicitly constructed by combining techniques from quantum error correction and renormalization-group flow Cong et al. (2019); Lake et al. (2022); Herrmann et al. (2022); Zapletal et al. (2024), which is a well-established method for the classification of phases of matter Sachdev (1999). Recently, QCNN s which are robust against errors in quantum data Zapletal et al. (2024), were implemented on a superconducting quantum processor under NISQ conditions Herrmann et al. (2022).
So far, QCNNs have, however, only been applied to one-dimensional systems, which can be accurately simulated on classical numerical computers. Their extension to two- and higher-dimensional systems, where quantum computers can potentially provide a computational advantage, has remained largely unexplored. Existing work includes a method for detecting topological phases in two-dimensions called Locally Error-corrected Decoration (LED) Cong et al. (2024), which is inspired by QCNNs. However, all current results are limited to snapshot-based LED, i.e. the classical post-processing of direct measurements. Thus, in its current form, it faces limitations once the characterization of the quantum states requires more data than can be efficiently processed on classical computers. This can occur at interesting points, as for example close to the critical point of a quantum phase transition. Hence extending QCNNs to two dimensions is a non-trivial generalization as quantum phases in two- and higher-dimensional systems are fundamentally different due to long-range entanglement, giving rise to topological order, which is absent in one dimension Wen (2017).
@@
Here, we address this open challenge by generalizing QCNNs to detect $\mathbb{Z}_{2}$-topological order in two-dimensional spin arrays. We analytically construct a QCNN inspired by quantum error correction and multiscale entanglement renormalization ansatz Vidal (2008). Using matrix-product-state simulations, we show that the QCNN recognizes the topological phase of the Toric Code Hamiltonian in a magnetic field from a topologically trivial phase. In contrast to existing methods Cong et al. (2024), which struggle to correctly classify complex quantum states close to phase boundaries, the QCNN precisely identifies the critical value of the magnetic field where the Toric Code undergoes the phase transition. Moreover, we show that the QCNN output is robust against incoherent errors below a threshold error probability allowing for the application of the QCNN on current and near-term quantum computers.
Classical approaches based on direct measurements, such as the snapshot-based LED method in Cong et al. (2024), cannot fully capture quantum correlations. In the case of the Toric Code, the stabilizer elements would require readout in different Pauli bases and thus cannot be simultaneously measured via direct measurements on typical quantum processors. In contrast, in processing quantum states via the QCNN quantum circuit, correlations between all stabilizer elements can be taken into account, increasing the threshold probability for errors that lead to such correlations, similarly to correlated quantum error correction Fowler (2013). As a result, the QCNN correctly identifies the topological phase in the presence of strong correlated errors, where classical approaches misclassify topological states. Note that, in contrast to quantum error correction constructions Krinner et al. (2022), it does not require additional readout qubits.
The QCNN we consider is a quantum circuit that can be written as a unitary
$U_{\mathrm{QCNN}}=U_{\mathrm{FC}}U_{\mathrm{P}}^{(d)}...U_{\mathrm{P}}^{(2)}U_{\mathrm{P}}^{(1)}U_{\mathrm{C}}\,$, (1)
consisting of a convolutional layer $U_{\mathrm{C}}$, $d$ pooling layers $U_{\mathrm{P}}^{(l)}$, for l=1,2,,dl=1,2,...,d, and a fully connected layer $U_{\mathrm{FC}}$ Herrmann et al. (2022); Zapletal et al. (2023a). The processing of quantum states in the QCNN is schematically depicted in Fig. 2. The convolutional layer $U_{\mathrm{C}}$ performs a translationally invariant transformation. In each pooling layer, the system size is reduced by discarding a fraction of qubits after applying gates on the remaining qubits. The fully connected layer $U_{\mathrm{FC}}$ prepares the output state of the remaining qubits, which are subsequently measured.
We construct the convolutional layer from Clifford gates and the pooling layers from single-qubit Pauli gates controlled by qubits that are subsequently discarded. In this case, the QCNN circuit Eq. 8 featuring only the single convolutional layer is equivalent to a deeper structure proposed in Ref. Cong et al. (2019), consisting of $d$ interleaved convolutional and pooling layers Herrmann et al. (2022); Zapletal et al. (2023a). The QCNN thus performs the same unitary with a substantially shorter circuit, see the supplemental material SM for details. Moreover, all such pooling layers can be efficiently performed in classical post-processing and only the initial convolutional layer $U_{\mathrm{C}}$ is executed on the quantum device.
@@
At each layer $l$, we can generate an output by ‘measuring’ the qubits in the computational basis,
$M^{l}_{\mathrm{QCNN}}=\dfrac{1}{N_{l}}\sum_{j=1}^{N_{l}}m_{j},\,$ (2)
where $m_{j}\in\{-1,1\}$ is the measurement outcome of qubit $j$ and $N_{l}$ the number of qubits in layer $l$. The measurement outcomes $m_{j}$ correspond to the values of plaquette and star operators on a lattice with the reduced system size $N_{l}$. The output $M^{l}_{\mathrm{QCNN}}$ thus measures how close the output state is to the error-free ground state of the Toric Code after having performed $l$ pooling layers.
As the pooling layers are performed in classical post-processing, their controlled Pauli gates are applied classically on the bit string of computational basis measurements for all qubits and the ’measurements’ correspond to simply recording the classical data. The support of the operator $M^{l}_{\mathrm{QCNN}}$ as defined in Eq. 2, will grow exponentially with the number of layers $l$. For our QCNN, we will obtain the following averages over a large number of samples
$\braket{\mathrm{QCNN}}\coloneqq\braket{M^{d}_{\mathrm{QCNN}}}=\begin{cases}1&\text{if topological phase}\\ 0&\text{if non-topological phase}\,.\end{cases}$ (3)
We focus on the two-dimensional Toric Code Kitaev (2006) with periodic boundary conditions. It is a natural candidate for exploring QCNNs for two-dimensional systems as it shows intrinsic $\mathbb{Z}_{2}$-topological order Kitaev (2006). The Toric Code is a two-dimensional square lattice with the qubits located on its edges. For exploring its condensed matter physics, it can be described by the Hamiltonian
$H_{\mathrm{TC}}=-\sum_{p}A_{p}-\sum_{v}B_{v},$ (4)
with the plaquette operators $A_{p}=\prod_{i=1}^{4}X_{i}$ and the star operators $B_{v}=\prod_{i=1}^{4}Z_{i}$, where $X_{i}$ ($Z_{i}$) is the Pauli-X(Z) operator on qubit $i$, see Fig. 2.
@@
Figure 2: QCNN structure. The first lattice on the input side shows the Toric Code lattice with dimension $d=4\times 4$, which is periodic in vertical and horizontal direction. The qubits on the respective sublattices are represented by gray (horizontal edges) and green circles (vertical edges). Examples of plaquette ($A_{p}$) and vertex ($B_{v}$) operators are highlighted in blue and orange. The circuit of the convolution maps the stabilizers of the input lattice to measurements on individual qubits of the respective colors, i.e., blue (green) qubits carry the measurement of the plaquette (vertex) stabilizer that is anchored on the qubits to their left. This convolution evolves error patterns on the input state and brings them into a form that can be corrected by the local operations in the pooling if the input was in the topological phase. Iterated pooling procedures reduce the lattice dimension (green squares) with every step until only two output qubits are left (green dots) and read out.
@@
Exposing the Toric Code to an external magnetic field gives rise to a rich phase diagram featuring the topologically ordered phase and a trivial phase, separated by first- or second-order phase transitions depending on the orientation of the field Trebst et al. (2007); Dusuel et al. (2011). To explore these regimes, we modify Eq. 4 to include applied magnetic fields in X/Y/Z-direction with strengths $h_{X}$, $h_{Z}$ and $h_{Y}$,
$H_{\mathrm{mag}}=H_{\mathrm{TC}}-h_{Z}\sum_{i}^{N}Z_{i}-h_{X}\sum_{i}^{N}X_{i}-h_{Y}\sum_{i}^{N}Y_{i}\,.$ (5)
@@
The QCNN that we apply to (noisy) ground states of this model is sketched in Fig. 2 and its detailed description can be found in the supplemental material SM . The QCNN we here propose is capable of detecting the topologically ordered phase with high accuracy. This means that the QCNN’s output converges to 1 with increasing depth $d$ if the input state is in the topological phase and to 0 otherwise.
@@
To evaluate its performance, we carried out several numerical simulations, where we use Matrix Product State (MPS) simulations Vidal (2007) to approximate ground states for large lattices Eq. 5 and classical simulations of Clifford gates on Pauli strings to explore the effects of incoherent Pauli noise. We show the technical details of our computations in the supplemental material SM .
@@
We first test the QCNN on ground states of $H_{\textrm{mag}}$ with a magnetic field in Z-direction, $h_{X}=h_{Y}=0$ in Eq. 5. Fig. 3 shows the resulting phase diagram, where we consider field strengths $0\leq h_{Z}\leq 1$. From iDMRG simulations for our cylindrical lattice, we determine the phase boundary at $h_{Z}^{*}=0.34$ as a peak in the second derivative $d^{2}E/dh_{Z}^{2}$ of the ground state energy with respect to $h_{Z}$, which we mark in the phase diagram. The plot of the QCNN output $\braket{M^{l}_{\mathrm{QCNN}}}$, see Eq. 2, shows a transition at this expected point up to the sample resolution. Additionally, the sharpness of the transition increases with the layer number in the sense that the absolute value of the gradient of $\braket{M^{l}_{\mathrm{QCNN}}}$ increases, which makes the transition in the output more step-like with an increasing number of pooling layers, thus showing the successful recognition of the topological phase by the QCNN.
@@
Figure 3: QCNN output for a variation of the magnetic field strength $h_{Z}$, where $h_{X}=0$, and for different depths $d$ of the QCNN, corresponding to the number of pooling layers $U_{\mathrm{P}}^{(l)}$. Each data point represents 2000 samples from the MPS representation of the corresponding ground state with the bond dimension $\chi=1250$ for an infinite cylinder with the periodic dimension $l_{2}=9$. The shaded region corresponds to the topologically ordered phase and the vertical line to critical magnetic field strength $h^{*}_{Z}=0.34$. The inset shows the output for different samples with higher resolution around the phase transition, which are calculated with a bond dimension of $\chi=1500$. For all samples, we observe increased steepness at the phase transition in the QCNN output with increasing QCNN depth corresponding to greater precision in the phase recognition.
@@
Furthermore, we test whether our QCNN is able to recognize the topological phase in the presence of incoherent noise. We do this by considering the Toric Code ground state perturbed by Pauli noise. Specifically, we apply random Pauli noise to each individual qubit in the form of the error channel